# Automated Transparency: A Legal and Empirical Analysis of the Digital Services Act Transparency Database

### v1

@@ Abstract.
Digital Services Act, transparency, computational compliance
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: the 2024 ACM Conference on Fairness, Accountability, and Transparency June 03–06, 2024Rio de Janeiro, Brazil; June 03–06, 2024, Rio de Janeiro, Brazilbooktitle: ;isbn: 978-1-4503-XXXX-X/18/06ccs: Social and professional topics Government technology policyccs: Information systems Web applicationsccs: Human-centered computing Collaborative and social computing design and evaluation methods
@@ 5.2. How Moderation Decisions and Detection are Done
Two attributes in the SoRs describe the manner of moderation: $automated\_detection$ and $automated\_decision$. The $automated\_detection$ field is binary and indicates whether automated tools were used to detect non-compliant content. Once content is detected as non-compliant, the decision on how to respond can be done in fully automated, partially automated, or in non automated manner, which is captured in $automated\_decision$ attribute in SoRs. Figure 2 presents the distribution of automated detection and decision in our dataset. We find that X never reports automated means for detection, some platforms (such as Pinterest and Google Play) do not use automated means for detection in half of their SoRs, and others (such as App Store, Google Shopping and TikTok) rely mostly on automated detection. Regarding the decision for content moderation, X, LinkedIn and Booking.com do not automate at all, Facebook, Instagram, Snapchat, and YouTube make automated decisions for some SoRs, whereas AliExpress, Google Maps, Google Shopping, and TikTok make fully automated decisions for most (or all) SoRs. Also, we find that automated detection and fully automated decision are highly correlated in the moderation process, as evident in Figure 2, where the height of bars corresponding to automated detection (blue) and automated decision (green) appear correlated. When platforms do not perform automated detection (orange), then they are more likely to perform partial automated decision rather than taking no automated decision or fully automated decision.
@@ 5.3. Why: Reasons for Moderation
Illegal content. For illegal content, two attributes ($illegal\_content\_legal\_ground$, that includes up to 500 characters and $illegal\_content\_explanation$, up to 2000 characters) are specified in the Transparency Database for platforms to provide details of illegal content. Table 1 gives the distribution of the most prominent grounds ($illegal\_content\_legal\_ground$) for marking content as illegal: AliExpress (66.2%) tops the list, followed by X (18.7%), Google Maps (8.3%), YouTube (4.3%), TikTok (2.0%), Google Play (1.1%), Facebook (0.07%), and Instagram (0.02%). Only 5 SoRs submitted by Google Shopping are flagged as illegal content. It appears that some VLOPs like Google Maps have mostly defamation related content whereas other VLOPs like TikTok have wide varying types of illegal content. We also notice that some VLOPs like Facebook specify only one standard reason  [^7] For example ‘This was a violation of sections 3.2 of our terms of service’. for reporting all illegal content. We also find that the optional attribute $decision\_ground\_reference\_url$ which is supposed to provide URL to the ToS or law, contains NULL values 99.9% of the times. Having said that, there are additional attributes in the Transparency Database which give more details about the illegal and ToS violation content. Further, we looked at values of $illegal\_content\_explanation$ which gives VLOPs more opportunity to give detailed explanation of grounds of flagging illegal content. To our surprise, only AliExpress makes use of this option. Out of 261,458 SoRs marked as illegal content, AliExpress gives specific and detailed explanations for 173,292 (66.2%) of them, mostly in Chinese, followed by Spanish and French. Besides AliExpress, YouTube (4%) also provides explanations with userID and product name anonymized.
ToS violations. Google Shopping (52.2%) submits more than half of the ToS violations followed by TikTok (17.1%), Amazon (10.8%), Facebook (9.8%), Pinterest (3.8%), Google Maps (2.6%), YouTube (1.8%), Instagram (0.7%), AliExpress (0.3%), and Google Play (0.2%). The trend is completely different in terms of illegal content submissions. For ToS violation content, attributes namely $incompatible\_content\_ground$ (upto 500 characters) and $incompatible\_content\_explanation$ (upto 2000 characters) are specified in the Database for VLOPs to use to provide details of ToS violation content. Table 2 gives distribution of most prominent grounds ($incompatible\_content\_ground$) for marking a ToS violation content by most active VLOPs submitting ToS violation content. As an illustration, Google Shopping does not provide any URLs to their ToS for the violations. However, they provide grounds for ToS violations, most of which is due to promotional overlays on images. Likewise, TikTok mentions harassment and bullying as the main grounds for ToS violations.
@@
In addition, the $decision\_facts$ attribute is mandatory for SoRs, and it is central given that it describes the facts and circumstances relied on in taking the content moderation decision, which can help provide information about how platforms contextualize the complex rules they are bound to apply at scale. However, we find that this attribute mostly features generic templates that may not provide very meaningful information. Google Shopping, the most active VLOPs, submits only a standard template888‘When reviewing content or accounts to determine whether they are illegal or violate our policies, we take various information into consideration when making a decision, including product data, website quality, merchant information, account information (e.g., past history of policy violations), and other information provided through reporting mechanisms (where applicable) and own-initiative reviews.’ in most (99.5%) of the SoRs. The next key contributor of SoRs, TikTok, uses a similar albeit shorter statement999‘The decision was taken pursuant to own-initiative investigations.’ in 98.5% of its SoRs. AliExpress seems to be the only platform providing more specific details around moderation. [^10] For instance, ‘The brand name of your product [id:abc] has been selected incorrectly and therefore cannot be published’; or ‘To xyz: Your product [id: abc] is similar to the def logo and important product elements, which can easily cause confusion among consumers, so it cannot be published’.
@@ 5.5. Where Content is Moderated: Language and Territorial Scope
Language is often a proxy indicator of the region from where the content was posted, or of a platform’s operational language in a given geographical area. In this section, we offer some language-based insights for SoRs, and we also revisit some earlier observations in the light of these insights. The optional attribute $content\_language$ in the SoR schema is only populated in approximately 50% of the records in our dataset. German (DE, 34.2%) is the most frequent language, followed by English (EN, 26.5%), French (FR, 9.9%), Italian (IT, 9.3%), Spanish (SP, 3.3%), Polish (PL, 2.9%), Dutch (NL, 2.7%), Czech (CS, 2%), Romanian (RO, 1.7%), and Slovak (SK, 1.4\$). Only three VLOPs, namely, Google Shopping, Google Play, and Booking.com, fill the language of content.
Additionally, SoRs feature several text fields that can be used as a more accurate proxy of the language of the moderated content. Overall, we inspect the language used in the following fields that require platforms to enter text in free form (with some limit in the number of characters): (a) Facts and circumstances relied on in taking the decision (decision_facts) (limit of 2000 characters, (b) Illegal Content Legal Grounds and Illegal Content Explanation are two fields that are required in a SoR if the DECISION_GROUND_ILLEGAL_CONTENT was the decision_ground, (c) Incompatible Content Grounds and Incompatible Content Explanation are required if DECISION_GROUND_INCOMPATIBLE_CONTENT was the decision_ground. For the decision_facts, all platforms with the exception of AliExpress (where the dominant language is Chinese) report content in English. Notably, there is a difference in the length of the description with the shortest facts being provided by Instagram and Facebook (51 and 55 characters respectively) and X (59 characters) and the longest facts being provided by Google Maps/Play/Shopping (438/468/433 characters respectively) and YouTube (635 characters). These numbers also correlate with the number of unique entries by platforms in this field, e.g. Instagram and Facebook provide only 2 entries here, namely ”This content was determined to violate terms of use” and ”This account was determined to violate terms of use”. For the Incompatible Content Grounds, all platforms except AliExpress and TikTok post content only in English (in total 13,0861,024 SoRs). For AliExpress the dominant language is Chinese (¿99% of cases), while TikTok provides grounds in 31 languages. More specifically, in the cases of booking.com, Facebook, and Instagram there is only one standard entry for all relevant SoRs: Violation of article 1F in section How we work in Terms and Condition, This was a violation of ”sections 3.2” of our terms of service and This was a violation of ”How You Can‘t Use Instagram section” respectively. Notably, for the Incompatible Content Explanation and AliExpress, there is some incompatibility between the language of the grounds (Chinese) and of the explanation (could be Spanish).
@@
A different location-based attribute in the SoRs is $territorial\_scope$, which records the territorial scope of the restrictions imposed - in other words, the geographical limitations applied to content visibility, which we could also refer as the geo-blocking of content as a moderation sanction. 72.68% of the SoRs apply territorial limitations to the entire European Economic Area (EEA), followed by 16.34% to the entire European Union. There are also a few specific national level restrictions applied to Germany, France, Spain, and Italy, respectively.

### v2

@@ Abstract.
Digital Services Act, Transparency, Computational Compliance.
journalyear: 2024copyright: rightsretainedconference: ACM Conference on Fairness, Accountability, and Transparency; June 3–6, 2024; Rio de Janeiro, Brazilbooktitle: ACM Conference on Fairness, Accountability, and Transparency (ACM FAccT ’24), June 3–6, 2024, Rio de Janeiro, Brazildoi: 10.1145/3630106.3658960isbn: 979-8-4007-0450-5/24/06ccs: Social and professional topics Government technology policyccs: Information systems Web applicationsccs: Human-centered computing Collaborative and social computing design and evaluation methodsccs: Applied computing Law, social and behavioral sciences
@@ 5.2. How Moderation Decisions and Detection are Done
Two attributes in the SoRs describe the manner of moderation: ‘automated detection’ and ‘automated decision’. The ‘automated detection’ field is binary and indicates whether automated tools were used to detect non-compliant content. Once content is detected as non-compliant, the decision on how to respond can be done in fully automated, partially automated, or in non automated manner, which is captured in ‘automated decision’ attribute in SoRs. Figure 2 presents the distribution of automated detection and decision in our dataset. We find that X never reports automated means for detection, some platforms (such as Pinterest and Google Play) do not use automated means for detection in half of their SoRs, and others (such as App Store, Google Shopping and TikTok) rely mostly on automated detection. Regarding the decision for content moderation, X, LinkedIn and Booking.com do not automate at all, Facebook, Instagram, Snapchat, and YouTube make automated decisions for some SoRs, whereas AliExpress, Google Maps, Google Shopping, and TikTok make fully automated decisions for most (or all) SoRs. Also, we find that automated detection and fully automated decision are highly correlated in the moderation process, as evident in Figure 2, where the height of bars corresponding to automated detection (blue) and automated decision (green) appear correlated. When platforms do not perform automated detection (orange), then they are more likely to perform partial automated decision rather than taking no automated decision or fully automated decision.
@@ 5.3. Why: Reasons for Moderation
Illegal content. For illegal content, two attributes (‘illegal content legal ground’, that includes up to 500 characters and ‘illegal content explanation’, up to 2000 characters) are specified in the Transparency Database for platforms to provide details of illegal content. Table 1 gives the distribution of the most prominent grounds (‘illegal content legal ground’) for marking content as illegal: AliExpress (66.2%) tops the list, followed by X (18.7%), Google Maps (8.3%), YouTube (4.3%), TikTok (2.0%), Google Play (1.1%), Facebook (0.07%), and Instagram (0.02%). Only 5 SoRs submitted by Google Shopping are flagged as illegal content. It appears that some VLOPs like Google Maps have mostly defamation related content whereas other VLOPs like TikTok have wide varying types of illegal content. We also notice that some VLOPs like Facebook specify only one standard reason  [^7] For example ‘This was a violation of sections 3.2 of our terms of service’. for reporting all illegal content. We also find that the optional attribute ‘decision ground reference url’ which is supposed to provide URL to the ToS or law, contains NULL values 99.9% of the times. Having said that, there are additional attributes in the Transparency Database which give more details about the illegal and ToS violation content. Further, we looked at values of ‘illegal content explanation’ which gives VLOPs more opportunity to give detailed explanation of grounds of flagging illegal content. To our surprise, only AliExpress makes use of this option. Out of 261,458 SoRs marked as illegal content, AliExpress gives specific and detailed explanations for 173,292 (66.2%) of them, mostly in Chinese, followed by Spanish and French. Besides AliExpress, YouTube (4%) also provides explanations with userID and product name anonymized.
ToS violations. Google Shopping (52.2%) submits more than half of the ToS violations followed by TikTok (17.1%), Amazon (10.8%), Facebook (9.8%), Pinterest (3.8%), Google Maps (2.6%), YouTube (1.8%), Instagram (0.7%), AliExpress (0.3%), and Google Play (0.2%). The trend is completely different in terms of illegal content submissions. For ToS violation content, attributes namely ‘incompatible content ground’ (upto 500 characters) and ‘incompatible content explanation’ (upto 2000 characters) are specified in the Database for VLOPs to use to provide details of ToS violation content. Table 2 gives distribution of most prominent grounds (‘incompatible content ground’) for marking a ToS violation content by most active VLOPs submitting ToS violation content. As an illustration, Google Shopping does not provide any URLs to their ToS for the violations. However, they provide grounds for ToS violations, most of which is due to promotional overlays on images. Likewise, TikTok mentions harassment and bullying as the main grounds for ToS violations.
@@
In addition, the ‘decision facts’ attribute is mandatory for SoRs, and it is central given that it describes the facts and circumstances relied on in taking the content moderation decision, which can help provide information about how platforms contextualize the complex rules they are bound to apply at scale. However, we find that this attribute mostly features generic templates that may not provide very meaningful information. Google Shopping, the most active VLOPs, submits only a standard template888‘When reviewing content or accounts to determine whether they are illegal or violate our policies, we take various information into consideration when making a decision, including product data, website quality, merchant information, account information (e.g., past history of policy violations), and other information provided through reporting mechanisms (where applicable) and own-initiative reviews.’ in most (99.5%) of the SoRs. The next key contributor of SoRs, TikTok, uses a similar albeit shorter statement999‘The decision was taken pursuant to own-initiative investigations.’ in 98.5% of its SoRs. AliExpress seems to be the only platform providing more specific details around moderation. [^10] For instance, ‘The brand name of your product [id:abc] has been selected incorrectly and therefore cannot be published’; or ‘To xyz: Your product [id: abc] is similar to the def logo and important product elements, which can easily cause confusion among consumers, so it cannot be published’.
@@ 5.5. Where Content is Moderated: Language and Territorial Scope
Language is often a proxy indicator of the region from where the content was posted, or of a platform’s operational language in a given geographical area. In this section, we offer some language-based insights for SoRs, and we also revisit some earlier observations in the light of these insights. The optional attribute ‘content language’ in the SoR schema is only populated in approximately 50% of the records in our dataset. German (DE, 34.2%) is the most frequent language, followed by English (EN, 26.5%), French (FR, 9.9%), Italian (IT, 9.3%), Spanish (SP, 3.3%), Polish (PL, 2.9%), Dutch (NL, 2.7%), Czech (CS, 2%), Romanian (RO, 1.7%), and Slovak (SK, 1.4\$). Only three VLOPs, namely, Google Shopping, Google Play, and Booking.com, fill the language of content.
Additionally, SoRs feature several text fields that can be used as a more accurate proxy of the language of the moderated content. Overall, we inspect the language used in the following fields that require platforms to enter text in free form (with some limit in the number of characters): (a) Facts and circumstances relied on in taking the decision (decision_facts) (limit of 2000 characters, (b) Illegal Content Legal Grounds and Illegal Content Explanation are two fields that are required in a SoR if ‘decision ground’ was illegal content, (c) Incompatible Content Grounds and Incompatible Content Explanation are required if ‘decision ground’ was incompatible content. For the decision_facts, all platforms with the exception of AliExpress (where the dominant language is Chinese) report content in English. Notably, there is a difference in the length of the description with the shortest facts being provided by Instagram and Facebook (51 and 55 characters respectively) and X (59 characters) and the longest facts being provided by Google Maps/Play/Shopping (438/468/433 characters respectively) and YouTube (635 characters). These numbers also correlate with the number of unique entries by platforms in this field, e.g. Instagram and Facebook provide only 2 entries here, namely ”This content was determined to violate terms of use” and ”This account was determined to violate terms of use”. For the Incompatible Content Grounds, all platforms except AliExpress and TikTok post content only in English (in total 13,0861,024 SoRs). For AliExpress the dominant language is Chinese (¿99% of cases), while TikTok provides grounds in 31 languages. More specifically, in the cases of booking.com, Facebook, and Instagram there is only one standard entry for all relevant SoRs: Violation of article 1F in section How we work in Terms and Condition, This was a violation of ”sections 3.2” of our terms of service and This was a violation of ”How You Can‘t Use Instagram section” respectively. Notably, for the Incompatible Content Explanation and AliExpress, there is some incompatibility between the language of the grounds (Chinese) and of the explanation (could be Spanish).
@@
A different location-based attribute in the SoRs is ‘territorial scope’, which records the territorial scope of the restrictions imposed - in other words, the geographical limitations applied to content visibility, which we could also refer as the geo-blocking of content as a moderation sanction. 72.68% of the SoRs apply territorial limitations to the entire European Economic Area (EEA), followed by 16.34% to the entire European Union. There are also a few specific national level restrictions applied to Germany, France, Spain, and Italy, respectively.
@@ 6. Conclusion: Does the DSA’s automated transparency rise to its promise?
Acknowledgements.
CG is supported by the ERC Starting Grant research project HUMANads (ERC-2021-StG No 101041824) and the Spinoza grant of the Dutch Research Council (NWO), awarded in 2021 to José van Dijck, Professor of Media and Digital Society at Utrecht University.